# MERba: Multi-Receptive Field MambaVision for Micro-Expression Recognition

### v1

@@ IV-D2 Asymmetric Multi-Scanning Strategy
We first investigate the impact of the number of scanning directions on MER performance. As shown in Table VII, compared with using a single ($a$) or a pair of directions ($a,b$), introducing more scanning directions leads to consistent performance gains across UF1, UAR, and ACC. This confirms that incorporating diverse scanning combinations enhances the spatial perception capability of the local extractor in ME feature extraction. However, we also observe that adding bi-directional ($a_{bi}$, $b_{bi}$) or horizontal symmetric ($a_{sy}$, $b_{sy}$) counterparts to the base directions does not lead to significant improvements. In contrast, utilizing our proposed asymmetric multi-scanning directions ($a,b,c,d$) achieves the best performance, indicating its ability to capture the most discriminative motion patterns without introducing unnecessary directional overlap.
@@ IV-D3 Dual-Granularity Classification Module
In Table II, we report the MER performance both with and without DGCM. It can be observed that introducing DGCM consistently improves performance on both DFME-public Test A and B, yielding UF1 gains of 2.31% and 1.88%, respectively. Furthermore, the confusion matrices in Figure 3 provide deeper insights into the impact of DGCM. Notably, the number of true positives for three $negative$ emotions—$disgust$, $fear$, and $sadness$—increases significantly. In addition, the tendency to misjudge $fear$ as $anger$/$disgust$, as well as the reciprocal confusion between $anger$ and $disgust$, is noticeably alleviated. These observations confirm DGCM’s effectiveness in disambiguating fine-grained negative emotions. However, this improvement comes with a trade-off: accuracy for certain non-negative emotions such as $surprise$ slightly decreases, possibly due to the model focusing more on subtle negative distinctions at the expense of broader category coverage.
@@ V Conclusion
In this paper, we proposed MERba, a novel architecture for MER, designed to address key challenges in feature extraction, global dependency modeling, and high inter-class similarity of MEs. By combining local extractors with global self-attention mechanisms, MERba effectively captures both fine-grained facial features and long-range dependencies, which are crucial for recognizing MEs. The asymmetric multi-scanning strategy reduces redundancy in scanning while enhancing the model’s spatial perception. Additionally, DGCM introduces a dual-head coarse-to-fine paradigm that disentangles emotion polarity from subtle category distinctions, enhancing recognition of highly similar MEs. The experimental results on benchmark MER datasets DFME-public and 3DB-Combined demonstrate that MERba outperforms existing methods, setting new state-of-the-art performance.

### v2

@@ IV-D2 Asymmetric Multi-Scanning Strategy
We first investigate the impact of the number of scanning directions on MER performance. As shown in Table VIII, compared with using a single ($a$) or a pair of directions ($a,b$), introducing more scanning directions leads to consistent performance gains across UF1, UAR, and ACC. This confirms that incorporating diverse scanning combinations enhances the spatial perception capability of the local extractor in ME feature extraction. However, we also observe that adding bi-directional ($a_{bi}$, $b_{bi}$) or horizontal symmetric ($a_{sy}$, $b_{sy}$) counterparts to the base directions does not lead to significant improvements. In contrast, utilizing our proposed asymmetric multi-scanning directions ($a,b,c,d$) achieves the best performance, indicating its ability to capture the most discriminative motion patterns without introducing unnecessary directional overlap.
@@ IV-D3 Dual-Granularity Classification Module
In Table V and VI, we report the MER performance both with and without DGCM. It can be observed that introducing DGCM consistently improves performance on both DFME-public Test A and B, yielding UF1 gains of 2.31% and 1.88%, respectively. Similarly, on the MMEW dataset, DGCM leads to a 2.6% increase in overall accuracy, further validating its effectiveness across datasets. To gain deeper insight into its impact, the confusion matrices in Figure 3 visualize how DGCM alters class-wise predictions. Notably, the number of true positives (TP) for three negativenegative emotions—disgustdisgust, fearfear, and sadnesssadness—increases significantly. In addition, the tendency to misjudge fearfear as angeranger/disgustdisgust, as well as the reciprocal confusion between angeranger and disgustdisgust, is noticeably alleviated. These observations confirm DGCM’s effectiveness in disambiguating fine-grained negative emotions. However, this improvement comes with a trade-off: accuracy for certain non-negative emotions such as surprisesurprise slightly decreases, possibly due to the model focusing more on subtle negative distinctions at the expense of broader category coverage.
IV-E Case Study
To better understand the inner workings of our model, we conduct a visual case study using Gradient-weighted Class Activation Mapping (Grad-CAM)[34]. As shown in Figure 4, we compare the attention heatmaps generated by MambaVision-B and our MERba-DGCM on three unseen test samples from the DFME dataset, covering the emotion categories of negative (disgust), positive, and surprise. For both models, Grad-CAM is applied to the last layer of the feature extractor to enable a direct comparison of high-level spatial attention.
A closer look reveals that MambaVision-B tends to spread attention over irrelevant areas, sometimes mistakenly focusing on background regions, including the subject’s hair or the image boundaries. In contrast, our MERba-DGCM consistently attends to semantically meaningful facial areas, such as the nasolabial folds for disgust, the cheek and lip corners for positive, and the widened eye region for surprise. These qualitative results complement our quantitative findings, confirming that MERba-DGCM captures more discriminative and interpretable motion cues through its integrated local-global modeling and coarse-to-fine classification strategy.
Figure 4: Grad-CAM visualizations of ME samples from DFME.
@@ V Conclusion
In this paper, we proposed MERba, a novel architecture for MER, designed to address key challenges in feature extraction, global dependency modeling, and high inter-class similarity of MEs. By combining local extractors with global self-attention mechanisms, MERba effectively captures both fine-grained facial features and long-range dependencies, which are crucial for recognizing MEs. The asymmetric multi-scanning strategy reduces redundancy in scanning while enhancing the model’s spatial perception. Additionally, DGCM introduces a dual-head coarse-to-fine paradigm that disentangles emotion polarity from subtle category distinctions, enhancing recognition of highly similar MEs. The experimental results on benchmark MER datasets (3DB-Combined, DFME-public, and MMEW) demonstrate that MERba outperforms existing methods, setting new SOTA performance.